### 19 Jan 2024 v6

# Towards domain-invariant Self-Supervised Learning with Batch Styles Standardization

Towards domain-invariant Self-Supervised Learning with Batch Styles Standardization
(Supplementary material)
Marin Scalbert & Maria Vakalopoulou
MICS
CentraleSupélec, Université Paris-Saclay
Gif-sur-Yvette, France
{marin.scalbert,maria.vakalopoulou}@centralesupelec.fr
&Florent Couzinié-Devy
VitaDX
Paris, France
f.couzinie-devy@vitadx.com
Appendix A Pseudo-code and PyTorch implementation of Batch Styles Standardization
On Algorithm 1 and Listing 1, a pseudo-code along with a PyTorch implementation of Batch Styles Standardization are provided. The full code will be released upon acceptance.
Input :
  • $\{{\bm{X}}_{i}\}_{1\leq i\leq N}$: Batch of images
  • $(r_{min},r_{max})$: Minimum and maximum area ratios between the substituted amplitudes components and the full amplitude.
Output :  Batch of images with standardized style: $\{\hat{{\bm{X}}}_{i}\}_{1\leq i\leq N}$
1 // Computes Fourier transform for all images in the batch
2 for $i\leftarrow 1$ to $N$ do
3       $\mathcal{A}({\bm{X}}_{i})\leftarrow\sqrt{\operatorname{Re}\left(\mathcal{F}({\bm{X}}_{i})\right)^{2}+\operatorname{Im}\left(\mathcal{F}({\bm{X}}_{i})\right)^{2}}$
4       $\mathcal{P}({\bm{X}}_{i})\leftarrow\arctan\left(\dfrac{\operatorname{Im}\left(\mathcal{F}({\bm{X}}_{i})\right)}{\operatorname{Re}\left(\mathcal{F}({\bm{X}}_{i})\right)}\right)$
5 end for
6// Sample the index $k$ associated to the style image and sample the amplitudes area ratio $r$
7 $k\sim U(\{1,\cdots,N\})$
8 $r\sim U(r_{min},r_{max})$
9 // For each image, substitute the low-frequency components with those of the style image then apply the inverse Fourier transform to transfer the style onto the original image.
10 for $i\leftarrow 1$ to $N$ do
11       $\hat{\mathcal{A}}({\bm{X}}_{i})\leftarrow\textnormal{{substitute\_low\_freq}}(\mathcal{A}({\bm{X}}_{i}),\mathcal{A}({\bm{X}}_{k}),r)$
12       $\hat{{\bm{X}}}_{i}\leftarrow\mathcal{F}^{-1}\left(\hat{\mathcal{A}}({\bm{X}}_{i})e^{-i\mathcal{P}({\bm{X}}_{i})}\right)$
13 end for
return $\{\hat{{\bm{X}}}_{i}\}_{1\leq i\leq N}$
Algorithm 1 Batch Styles Standardization
{listing*}
[ht]
class BatchStylesStandardization():
”””Implements Batch Styles Standardization. Given a
batch of N images and their Fourier transforms,
we manipulate the different amplitudes by
substituting their low-frequency components
with those a randomly chosen image.
[…]
ratios (tuple): \$(r_min, r_max)\$ specifying
the minimum and maximum possible
areas ratio between the substituted
amplitude and the full amplitude.
def __init__(self, ratios):
self.ratios = ratios
[…]
”””Substitute the low-frequency components
of the source amplitudes with those
of the target amplitudes.
[…]
src_amp (torch.Tensor): Source amplitudes
tgt_amp (torch.Tensor): Target amplitudes
ratio (float): Area ratio between the
substituted amplitude and the full
amplitude.
[…]
torch.Tensor: Source amplitudes where
the low-frequency components have been
substituted with those
%**** iclr2024_conference_supp.tex Line 175 ****of the target amplitudes.
# Compute center coordinates of amplitudes
h, w = src_amp.shape[-2:]
hc, wc = int(h//2), int(w//2)
[…]
# to be substituted
l = min([int(ratio*h/2), int(ratio*w/2)])
[…]
# amplitudes with those of the target amplitudes
low_freq_tgt_amp = tgt_amp[
…, hc-l:hc+l, wc-l:wc+l]
src_amp[
…, hc-l:hc+l, wc-l:wc+l] = low_freq_tgt_amp
return src_amp
[…]
”””Apply batch styles standardization n_views times
on a batch of \$N\$ images.
[…]
imgs (torch.Tensor): Batch of images (N, 3, H, W)
%**** iclr2024_conference_supp.tex Line 200 ****n_views (int): Number of augmented views
[…]
torch.Tensor: Batch with standardized styles
(N, n_views, 3, H, W)
# Apply FFT on source images
fft = torch.fft.fftn(
imgs, dim=(-2, -1))
# Shift low-frequency components to the center
fft = torch.fft.fftshift(
fft, dim=(-2, -1))
# Retrieve amplitude and phase
amp, phase = fft.abs(), fft.angle()
[…]
# ref styles
bs = imgs.size(0)
sampled_ind = torch.randperm(bs)[:n_views]
[…]
# of the n_views sampled images
src_amp = amp.unsqueeze(1).repeat(
[1, n_views, 1, 1, 1])
tgt_amp = amp[sampled_ind].unsqueeze(0).expand(
%**** iclr2024_conference_supp.tex Line 225 ****bs, -1, -1, -1, -1)
sampled_ratio = random.uniform(*self.ratios)
amp = self.substitute_low_freq(
src_amp, tgt_amp, sampled_ratio)
[…]
# Reconstruct FFT from amp and phase
fft = torch.polar(amp, phase)
# Shift back low-frequency to their
# original positions
fft = torch.fft.ifftshift(fft, dim=(-2, -1))
# Invert FFT
imgs = torch.fft.ifftn(
fft, dim=(-2, -1)).real.clamp(0, 1)
return imgs
Batch Styles Standardization PyTorch implementation
Appendix B Technical details about SSL methods
B.1 SimCLR
SimCLR aims to bring representations of augmented views of the same image closer (positives) while repelling all other images representations (negatives). In practice, given a batch of $N$ images, each image is augmented $V$ times independently resulting in a $N\times V$ images grid where each row $c$ corresponds to a content and each column $s$ to a view. For each image ${\bm{X}}_{cs}$ and its corresponding representation ${\bm{z}}_{cs}\in\mathbb{R}^{D}$, SimCLR minimizes the NT-Xent loss with temperature $T$:
$\mathcal{L}_{{\color[rgb]{0,0,0}c}{\color[rgb]{0,0,0}s}}=\displaystyle\dfrac{-1}{V-1}\sum_{s^{\prime}\neq{\color[rgb]{0,0,0}s}}\log\left(\dfrac{e^{{\bm{z}}_{{\color[rgb]{0,0,0}c}{\color[rgb]{0,0,0}s}}\cdot{\bm{z}}_{{\color[rgb]{0,0,0}c}s^{\prime}}/T}}{\displaystyle\sum_{(c^{\prime\prime},s^{\prime\prime})\neq({\color[rgb]{0,0,0}c},{\color[rgb]{0,0,0}s})}e^{{\bm{z}}_{{\color[rgb]{0,0,0}c}{\color[rgb]{0,0,0}s}}\cdot{\bm{z}}_{c^{\prime\prime}s^{\prime\prime}}/T}}\right)$ (1)
B.2 SWaV
SWaV computes the representations of different views of the same image while clustering them using an online algorithm. Since representations should capture similar information, SWaV assumes that one view’s cluster assignment can predicted from representations of other views. This swapped prediction idea is the core concept behind SWaV loss formulation.
Concretely, in SWaV, an image ${\bm{X}}_{n}$ is augmented into $2$ views ${\bm{X}}_{n}^{(s)}$ and ${\bm{X}}_{n}^{(t)}$, with corresponding representations ${\bm{z}}_{n}^{(s)}$ and ${\bm{z}}_{n}^{(t)}$. Similarities between representations and $K$ learnable cluster centroids/prototypes ${\bm{C}}\in\mathbb{R}^{K\times D}$ are computed and converted into probabilities such as follows:
${\bm{p}}_{n}^{(v)}$ $=\text{softmax}\left(\dfrac{{\bm{z}}_{n}^{(v)}\cdot{\bm{C}}^{T}}{\tau}\right)\ \ \forall v\in\{s,t\}$ (2)
To compute cluster assignments also referred to as codes and denoted ${\bm{q}}_{n}^{(v)}$, SWaV relies on the Sinkhorn-Klopp (SK) algorithm (cuturi2013sinkhorn). SK is performed on all views representations trying to assign representations to the most similar centroids but also uniformly among clusters. Finally, based on the swapped prediction concept, SWaV minimizes the following per-sample loss:
$\mathcal{L}_{n}=H({\bm{q}}_{n}^{(s)},{\bm{p}}_{n}^{(t)})+H({\bm{q}}_{n}^{(t)},{\bm{p}}_{n}^{(s)})$ (3)
In Equation 3, $H({\bm{p}},{\bm{q}})$ stands for the cross-entropy between an approximated probability distribution ${\bm{q}}$ and a true probability distribution ${\bm{q}}$.
In practice, SWaV employs a multi-crop strategy, generating $2$ global views (large crops) and $V$ local views (small crops) for each image. Cluster assignments are then computed only from the $2$ global views while probabilities derived from all the $V+2$ views. In this setting, SWaV minimizes the following loss:
$\mathcal{L}_{n}=\dfrac{1}{2(V+1)}\displaystyle\sum_{i=1}^{2}\sum_{v=1}^{V+2}\bm{1}_{i\neq v}H({\bm{q}}_{n}^{(i)},{\bm{p}}_{n}^{(v)})$ (4)
B.3 MSN
Given two views of the same image, MSN randomly masks the patches of one view and leaves the other unchanged. Then, MSN’s goal is to match the representation of the masked view with that of the unmasked view.
To derive a view’s representation, MSN computes the similarities between its embedding and a set of cluster centroids/prototypes, subsequently transforming them into a probability distribution. As direct matching of these representations can lead to representation collapse, MSN simultaneously optimizes a cross-entropy term along with an entropy regularization term on the mean representation of the masked views. The entropy regularization term encourages the model to use the entire set of centroids/prototypes. Additionally, MSN employs Sinkhorn-Klopp on the representations of the unmasked views to avoid tuning the hyperparameter weighting the entropy regularization term.
In practice and more formally, MSN generates for each image ${\bm{X}}_{n}$, $M$ masked views $\{{\bm{X}}_{n,1},\ldots,{\bm{X}}_{n,M}\}$ and a single unmasked view ${\bm{X}}_{n}^{+}$. Masked views are processed by a student encoder and the unmasked view by teacher encoder whose weights are updated via exponential moving average of the student encoder’s weights. Masked and unmasked views’ embeddings denoted $\{{\bm{z}}_{n,1},\ldots,{\bm{z}}_{n,M}\},{\bm{z}}_{n}^{+}$ are then compared to a set of centroids/prototypes ${\bm{C}}\in\mathbb{R}^{K\times D}$ and the resulting similarities are converted into probability distributions $\{{\bm{p}}_{n,1},\ldots,{\bm{p}}_{n,M}\},{\bm{p}}_{n}^{+}$:
$\begin{cases}{\bm{p}}_{n,m}=\text{softmax}\left(\dfrac{{\bm{z}}_{n,m}\cdot{\bm{C}}^{T}}{\tau}\right)\\ {\bm{p}}_{n}^{+}=\text{softmax}\left(\dfrac{{\bm{z}}_{n}^{+}\cdot{\bm{C}}^{T}}{\tau^{+}}\right)\end{cases}$ (5)
$\tau$ and $\tau^{+}$ stand for temperature hyperparameters and are chosen such that $\tau>\tau^{+}$ to encourage sharper probability distributions implicitly guiding the model to produce confident masked views representations. Given a batch of $N$ images, MSN minimizes the following loss:
$\begin{cases}\mathcal{L}=\dfrac{1}{NM}\displaystyle\sum_{n=1}^{N}\sum_{m=1}^{M}H({\bm{p}}_{n}^{+},{\bm{p}}_{n,m})-\lambda H(\bar{{\bm{p}}})\\ \bar{{\bm{p}}}=\dfrac{1}{NM}\displaystyle\sum_{n=1}^{N}\sum_{m=1}^{M}{\bm{p}}_{n,m}\end{cases}$ (6)
$H({\bm{p}},{\bm{q}})$ stands for the cross-entropy between an approximated probability distribution ${\bm{q}}$ and a true probability distribution ${\bm{q}}$ while $H(\bar{{\bm{p}}})$ denotes the entropy of the masked views’ mean representation $\bar{{\bm{p}}}$.
Appendix C Implementation details
C.1 Pretraining
On PACS and DomainNet, as part of the geometric augmentations, we use random crop resizing, horizontal flips, small rotations, cutout(devries2017improved) while color augmentations are applied in batch-wise manner using color jitter, random equalize, random posterize, random solarize and random grayscale. On Camelyon17 WILDS, we use random crop resizing, flips, rotations, cutout and batch-wise color jitter. All other hyperparameters for SimCLR, SWaV, MSN are respectively specified on Tables 1,  2,  3.
Table 1: Hyperparameters used for SimCLR extension based on Batch Styles Standardization
datasets PACS DomainNet Camelyon17 WILDS
backbone ResNet-18 ResNet-18 ResNet-50
$V$ $2\times 224^{2}+6\times 128^{2}$ $2\times 224^{2}+6\times 128^{2}$ $8\times 128^{2}$
$(r_{min},r_{max})$ $(0.02,1)$ $(0.02,1)$ $(0.02,0.1)$
$D$ $128$ $128$ $128$
$T$ $0.5$ $0.5$ $0.5$
$N$ $256$ $512$ $256$
steps $60$K $60$K $150$K
optimizer LARS LARS LARS
learning rate $0.2$ $0.4$ $0.2$
learning rate schedule linear warmp-up + cosine decay linear warmp-up + cosine decay linear warmp-up + cosine decay
weight decay $10^{-6}$ $10^{-6}$ $10^{-6}$
Table 2: Hyperparameters used for SWaV extension based on Batch Styles Standardization
datasets PACS DomainNet Camelyon17 WILDS
backbone ResNet-18 ResNet-18 ResNet-50
global views $2\times 224^{2}$ $2\times 224^{2}$ $2\times 128^{2}$
local views $6\times 128^{2}$ $6\times 128^{2}$ $6\times 128^{2}$
$(r_{min},r_{max})$ $(0.02,1)$ $(0.02,1)$ $(0.02,0.1)$
$K$ $256$ $256$ $256$
$D$ $128$ $128$ $128$
$\tau$ $0.1$ $0.1$ $0.1$
$N$ $256$ $512$ $256$
steps $60$K $60$K $150$K
optimizer LARS LARS LARS
learning rate $0.2$ $0.4$ $0.2$
learning rate schedule linear warmp-up + cosine decay linear warmp-up + cosine decay linear warmp-up + cosine decay
weight decay $10^{-6}$ $10^{-6}$ $10^{-6}$
Table 3: Hyperparameters used for MSN extension based on Batch Styles Standardization
backbone ViT-S/8
unmasked views $2\times 96^{2}$
masked views $10\times 64^{2}$
$(r_{min},r_{max})$ $(0.02,0.1)$
patch masking ratio $0.3$
$K$ $128$
$D$ $384$
$\tau$ $0.1$
$\tau^{+}$ $0.025$
$N$ $256$
steps $150$K
Optimizer LARS
learning rate $0.2$
learning rate schedule linear warm-up + cosine decay
weight decay $10^{-6}$
EMA momentum $0.995$
C.2 fine-tuning/Linear-probing
For all datasets (PACS, DomainNet, and Camelyon17 WILDS), we use the Adam optimization method (kingma2014adam) with an initial learning rate of $10^{-4}$, a learning rate scheduler with cosine decay, and weight decay of $10^{-4}$. The networks are trained respectively for $5$K, $1$K, and $15$K steps with batch sizes of $128$, $64$, and $64$. When performing linear probing, we follow the same normalization scheme as (he2022masked) by adding a batch normalization layer (ioffe2015batch) without affine parameters before the linear classifier.
Appendix D Discussions & additional visualizations
D.1 Transfer learning in DG/UDG
The usage of Transfer Learning in DG/UDG is common but we think it is misguided. Often the pretraining dataset, such as Imagenet, can include one or more of the target domains, e.g., photo for PACS or real for DomainNet. When evaluating on these domains, it is not possible to know if performances result from the generalization ability of the DG/UDG methods or from the transfer learning. For new DG/UDG methods, it is hard not to follow the common practice because transfer learning unfairly boosts the results of previous works, and state-of-the-art performances are often seen as a prerequisite for paper acceptance. We tried to limit the usage of transfer learning in our experiments and only used it for the DomainNet dataset. The bias introduced by Imagenet transfer learning can be seen in our results and especially for the domains that are close to Imagenet: on PACS (Table 1 main paper), without transfer learning, SimCLR with BSS is consistently the best method for all settings except for the photo domain, where methods using Imagenet transfer learning report better performance. However, on DomainNet (Table 2 main paper), when using Imagenet transfer learning similar to the other UDG methods, SimCLR with BSS achieves better performances on the real domain for 2 out of 3 proportions of labeled data ($1\%$, $5\%$). In the $10\%$ labeled data setting, where SimCLR with BSS is outperformed, it is worth noting that DiMAE is the only method to perform full fine-tuning which is probably the reason behind the gap of performances.
D.2 Features visualization
To assess the quality of the SSL representations and their ability to generalize across domains, we display, in Figure 1, t-SNE plots of the backbone representations for SimCLR with BSS and competitors on Camelyon17 WILDS.
Figure 1: t-SNE plots of the backbone representations for different UDG methods on Camelyon17 WILDS. Colors and markers correspond respectively to different domains and classes. On the target domains (hospital_1, hospital_2), our method (SimCLR w/ BSS) shows better domain confusion while keeping better class separability. Zoom on pdf for better visualization.
DARLING representations tend to be domain-invariant as lots of examples from different domains are superimposed. However, this is also the case for many examples from different classes indicating potentially poor model generalization. In contrast, DiMAE representations appear to be well separated by classes but also by domains, especially for the target domains hospital_1 and hospital_2, indicating a lack of domain-invariance. Finally, better class separability and domain confusion emerge from the representations of SimCLR with BSS revealing a better domain-invariance and a potentially better cross-domain generalization.
D.3 Impact of $r$ on BSS generated images
To illustrate the effect of the hyperparamter $r$ on the generated images by BSS, we apply BSS on a single batch fixing the chosen style image and varying $r$. The resulting images are reported in Figure 2. We can observe that as $r$ increases, textures/styles with higher frequencies are transferred to the resulting images.
Figure 2: Impact of hyperparameter $r$ on augmented images with BSS